OpenStudy (anonymous):

I'd be very grateful if someone could start me off on this: Express exp( (2x) / π ) cos (3x) as the real part of a single complex exponential. Hence find ∫ exp( (2x) / π ) cos 3x dx . (the limits I can't draw but are between 0 and π) @Physics

OpenStudy (anonymous):

I think it wants cos (3x) rewriting as an exponential.

OpenStudy (anonymous):

\[1/2 (e^{-3 i x}+ e^{3 i x})\]

OpenStudy (anonymous):

yes, then that turns out to be \[1/2(2e^{2x/\pi}+e^{-3 i x}+e^{3 i x })\] which is obviously equal to\[\cos(3x) + \sinh (2x/ \pi) + \cosh (2x/\pi)\]Which would be dead easy to integrate. Do you think that's right?

OpenStudy (anonymous):

So that gives as the final answer\[(\pi ^2 -4 e^2)/(2 \pi)\]

OpenStudy (anonymous):

integrating gave \[\left[ -3x \sin (3x) + 2/\pi \cosh (2x/ \pi) +2/\pi \sinh(2x/\pi) \right]_{0}^{\pi}\]

OpenStudy (jamesj):

What being asked here is this: what is the exponent f(x) such that Re [ exp(f(x)) ] = exp( (2x) / π ) cos (3x) And is clearly f(x) = ( 2/ π + 3i) x

OpenStudy (jamesj):

Hence the integral of exp( (2x) / π ) cos (3x) is the integral of the real part of exp( ( 2/ π + 3i) x ) which will turn out to be the real part of the integral of the same. I.e., \[ Re \left[ \int_0^{\pi} e^{(2/\pi + 3i)x} dx \right] \]

OpenStudy (anonymous):

let cos3x be = coskx coskx= (1/2i).[(exp(ikx)+exp(-ikx)] put this into the integral S exp(2x/pi). (1/2i).[(exp(ikx)+exp(-ikx)].dx (1/2i) is not the variable so goes out. (1/2i) S exp(2x/pi).[(exp(ikx)+exp(-ikx)].dx (1/2i) S [exp(2x/pi).(exp(ikx)+exp(2x/pi)exp(-ikx)].dx after thıs its quite easy just integral of an expo

OpenStudy (jamesj):

The question asks: "Express exp( (2x) / π ) cos (3x) as the real part of a single complex exponential." That's what I've done.

OpenStudy (anonymous):

I still don't get it. Which bit is the real part of the complex exponential and which part is the complex exponential?

OpenStudy (jamesj):

exp( (2x) / π ) cos (3x) = Re [ exp( (2x) / π ) . exp(3xi) ] = Re [ exp( (2/π+3i)x ) ]

OpenStudy (anonymous):

Oh right. I get it now. so you're saying exp( (2x) / π ) cos (3x) is only some of the truth. There's a complex bit which we've left out. The full thing is: exp( (2x) / π ) . exp(3xi) (the complex exponential) which can be integrated easily. Thank you so much.

OpenStudy (jamesj):

yes.

OpenStudy (anonymous):

If I could give you Graham's number of medals, I would.

OpenStudy (jamesj):

One will suffice ;-)